OpenStudy (anonymous):

solve (-9x+5)1/2=(3+4x)1/2

OpenStudy (anonymous):

\[\frac{ (-9x+5)1 }{ 2 }=\frac{ (3+4x)1 }{2 }\]

OpenStudy (anonymous):

2/13